OpenStudy (anonymous):

How do you graph??? -x + 4 What does Y =? Please help

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

im guessing y=-x+4

OpenStudy (anonymous):

how do I graph that?

OpenStudy (dumbcow):

to graph this line start at y_intercept (0,4) use slope to plot points from there slope = -1 (slope is always coefficient of x when in slope-intercept form) slope = rise/run go down 1 then over 1

OpenStudy (anonymous):

ok cool! that helps... now one more. what about... Y=3x - 2?

OpenStudy (dumbcow):

ok whats the y-intercept? what is the slope?

OpenStudy (dumbcow):

anything? y=mx+b m=slope b=y_intercept

OpenStudy (anonymous):

OMG I feel really retarded!

OpenStudy (dumbcow):

dont be :) ok slope is always number next to x slope = 3 y_intercept is always the constant y_intercept = -2 so start at (0,-2) for 1st point then use slope (up 3 over 1 )to get next point

OpenStudy (anonymous):

Holy crap I think I got that one! are u up for just 1 more?

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

2/5 (10x + 25) = -10

OpenStudy (dumbcow):

are we solving for x or graphing? where is y?

OpenStudy (anonymous):

Solving for X and there is no Y

OpenStudy (dumbcow):

ok distribute first 4x +10 = -10 subtract 10 on both sides 4x = -20 divide by 4 on both sides x = -5

OpenStudy (anonymous):

thank you very very much.i hope you can help me again next time... dumbcow...:)

OpenStudy (dumbcow):

:)